# FalconFS: Distributed File System for Large-Scale Deep Learning Pipeline

### v1

@@ 4.2.1 Hybrid Indexing Methods
When the coordinator detects load imbalance using statistics reported by MNodes, it conducts the load balancing algorithm as follows.
@@
    Redistribute $F$. The coordinator selects from the two redirection methods to approach load balancing. If we use path-walk redirection, assuming it evenly distributes all files with filename $F$ to all nodes, the inode numbers of $N_{max}$ and $N_{min}$ will be $\langle{N_{max}}\rangle-\frac{n-1}{n}|F|$ and $\langle{N_{min}}\rangle+\frac{1}{n}|F|$, respectively, where $|F|$ represents the number of files named $F$ in node $N_{max}$. If using overriding redirection for $F$, we will transfer all $|F|$ files from $N_{max}$ to $N_{min}$, yielding the inode numbers of $\langle{N_{max}}\rangle-|F|$ and $\langle{N_{min}}\rangle+|F|$, respectively. The coordinator chooses the method that minimizes the maximum inode count. After the method is chosen, the redirection entry is inserted to the exception table and the corresponding files are migrated among nodes. To ensure metadata consistency, access to the corresponding inodes are temporarily blocked during the migration.
@@ 4.3 Lazy Namespace Replication
Figure 7: Namespace Synchronization. Blue lines indicate remote lookup and read lines indicate invalidate.
Lazy namespace replication is another key enabler of the stateless-client architecture. FalconFS maintains a consistent but not necessarily complete namespace replica on each MNode and the coordinator, enabling local path resolution. To reduce overhead of maintaining consistency across all replicas, modifications to the namespace are lazily synchronized and an invalidation-based mechanism is adopted, inspired by cache coherence protocols and Hermes [20]. The design is guided by two principles:
@@
We will then introduce how the namespace replicas are maintained during related operations including creating/remove a directory, changing the permissions, and renaming.
@@
Figure 7 shows an example of how to create the directory /b. The client calculates the location of /b via the hybrid metadata indexing, and then sends a mkdir request to the corresponding MNode, i.e., $MNode_{0}$. Upon receiving the request, $MNode_{0}$ resolves the path by querying its local namespace replica to check the path existence and permissions. It also checks its inode table to confirm /b does not already exist. Once all checks are passed, $MNode_{0}$ creates an inode for /b in the inode table, adds a dentry for /b to its local namespace replica, and responds to the client.
Note that to retain the efficient single-hop processing of mkdir, $MNode_{0}$ does not proactively broadcast the new dentry to namespace replicas on other MNodes. Instead, the dentry information is fetched by other namespace replicas on demand. Specifically, when a MNode queries its local namespace replica for path resolution and finds a missing dentry, it fetches the information from the owner MNode calculated via the hybrid metadata indexing.
Figure 7 shows an example. When $MNode_{1}$ resolves the path /b/c, it finds the dentry for /b missing in its local namespace replica. It then sends a lookup request to $MNode_{0}$ (i.e., the owner MNode) to fetch the missing dentry to complete the path resolution and continue subsequent processing.
@@
In the example shown in Figure 7, the client sends an rmdir(/b) request to the coordinator. The coordinator acquires shared locks on all ancestor directories and an exclusive lock on the target directory to ensure path validity during execution. It then forwards the request to the directory inode’s owner, i.e., $MNode_{1}$ in the figure. $MNode_{1}$ locks /b’s inode to block subsequent lookup requests and broadcasts an invalidation request to other MNodes. Upon receipt, each MNode invalidates its local dentry of the target directory if it exists in the local namespace replica. Each MNode searches its inode table for entries whose key’s pid equals to /b’s ID (i.e., /b’s children), and responds to $MNode_{1}$ the existence of any children. $MNode_{1}$ aggregates these responses. If /b has no children, $MNode_{1}$ deletes the inode and notifies the coordinator, which will release locks and respond to the client. Otherwise, $MNode_{1}$ returns -ENOTEMPTY to abort the rmdir.
@@
FalconFS leverages namespace replicas to coordinate concurrent file requests. When a server (the coordinator or an MNode) processes a file request, it first resolves the path component by component in its local namespace replica. Dentry locks are acquired during the resolution. To improve parallelism, the coordinator acquires shared locks on intermediate dentries and an exclusive lock on the last component, while MNodes acquire shared locks on all dentries. If a dentry is missing in the local namespace replica, the server locks the dentry after it is retrieved from its owner MNode. Thus, concurrent requests on the same server are serialized by these locks. Concurrent requests on two different MNodes will not incur data race and thus can be executed in parallel. The last case is a request being processed on the coordinator ($Req_{C}$) and a request on an MNode ($Req_{M}$).
@@
The second case is that the invalidation is processed before $Req_{M}$ locks /a/b. In this case, $MNode_{M}$ will find the /a/b dentry invalid during path resolution of $Req_{M}$. $MNode_{M}$ then sends a lookup request to retrieve the dentry from $MNode_{C}$, who owns the /a/b inode. On $MNode_{C}$, the lookup will acquire a shared lock of the /a/b inode, which has already been locked by $Req_{C}$. Thus $Req_{M}$ will be blocked until $Req_{C}$ completes, forming a correct serialization of the two requests. Note that when processing the invalidation request, $MNode_{M}$ will discard all lookup responses whose requests are issued before the invalidation is received.
@@ Discussion.
Lazy namespace replication introduce one remote lookup for access to non-existent path. However, in DL workloads, such negative access is rare.
@@ 4.4 Concurrent Request Merging
Figure 8 illustrates the metadata servers’ request-handling mechanism, which employs concurrent request merging. Each MNode initializes a fixed number of database worker threads to serve as the backend for metadata storage and prepares a connection pool. The connection pool accepts incoming requests and puts them into request queues according to the request type. An idle worker thread retrieves a queue and executes all requests in the queue in a single database transaction, with the following optimizations.
@@
Due to the tree-structured nature of the file system namespace, requests can share common near-root path prefixes. As the request queue accumulates multiple operations, the worker coalesces shared path prefixes and eliminates redundant lock acquisitions. In Figure 8, the three create operations each walks two directories and one file. Rather than acquiring nine locks separately, the worker eliminates redundant lock acquisitions and acquires only six locks instead.
@@
Operations such as mkdir, create, and close modifies the inode table. To maintain file system metadata consistency, DFS typically warp each operation into a separate transaction to persist in atomic [33, 46, 27]. When an transaction commits, it synchronously appends the write-ahead log, leading to small writes that are unfriendly for storage. In FalconFS, as concurrent operations are batched into a single transaction, the worker coalesces small log appends into larger ones, improving the storage efficiency.
@@ 5 Implementation of VFS Compatibility
We observe that existing semantics provided to the lookup() method is sufficient to distinguish lookup intentions. Since Linux kernel 5.7, the VFS sets the global state flag LOOKUP_PARENT during path walk to indicate that the final component has not yet been reached —- a feature designed initially for the kernel audit subsystem [6], and the flag is passed to the lookup() method. If the flag is set, the client module knows that the lookup is for an intermediate directory and returns fake attributes (e.g., $mode=0777$, along with special $uid$ and $gid$ values) to pass VFS checking.
@@
Figure 9 illustrates an example of VFS shortcut. During a getattr operation for the path /a/b, the VFS resolves each component sequentially. It first looks up /, which results in a dcache hit. The VFS invokes the d_revalidate() method to validate the entry, receiving a positive response. Then, the VFS looks up /a, which misses the dcache. The lookup() method is called with the LOOKUP_PARENT flag set, returning a fake attribute. Finally, the VFS looks up /a/b, which also misses the dcache. Here, the lookup() method is invoked without flags. The module then issues a remote lookup request with the full path (/a/b) to the MNode, which executes the real path checking, executes the lookup operation, and returns the result to the module. The module returns the result to the VFS, completing the getattr operation.
@@ 6 Evaluation
Figure 11: Latency of metadata operations.
@@
    FalconFS’s performance is robust under adverse conditions like client memory limitations (§6.4) and load bursts (§6.5).
@@
    FalconFS’s performance in real-world DL workloads (§6.8).
@@ 6.1 Environment Setup
Testbed. We conduct experiments on a cluster of 13 dual-socket machines (configurations detailed in Table 2). To expand the test scale, we abstract each machine into two independent nodes, with each node bound to one socket, one SSD, and one NIC, scaling the testbed to 26 nodes. We restrict server resources to 4 cores per node to ensure clients can saturate the servers’s capability.
@@
We compare FalconFS with CephFS 12.2.13 [48], JuiceFS 1.2.1 [3] and Lustre 2.15.6 [4]. CephFS is a widely deployed DFS in data centers. JuiceFS is an open-source DFS targeting AI and data analytics workloads, and we deploy it with TiKV 1.16.1 [8] as its metadata engine and data storage. Lustre is a high-performance DFS widely used in HPC and data centers. CephFS is accessed via the libcephfs library due to observed instability in performance when using a VFS mount point. JuiceFS and Lustre are accessed through VFS mount points, as is FalconFS unless stated otherwise. FalconFS utilizes a modified FUSE kernel module and library that incorporate the optimizations described in §5. All DFSs disable metadata and data replication.
@@ 6.2 Metadata Performance
We scale the number of metadata servers from 4 to 16 and measure the peak throughput achievable by each file system. To saturate the metadata servers’ capacity, we gradually increase the number of client threads until the throughput no longer increases. When mounting with FUSE clients, we observe that 13 client nodes are insufficient to saturate FalconFS, so we present FalconFS’s performance using the LibFS interface, enabling each client node to generate higher concurrency. We ensure that the FUSE client and the LibFS client generate identical requests to the metadata servers; thus, given sufficient client nodes, FUSE clients would achieve comparable performance. In production environments, client nodes typically outnumber metadata servers and can fully saturate them. Figure 10 shows the results.
For the operations create and unlink, FalconFS achieves speedups of 0.82–2.26$\times$ on Lustre and higher gains on CephFS and JuiceFS. This performance improvement stems from two factors: (a) FalconFS do not maintain directory’s atime and mtime, eliminating the need to update directory metadata; (b) concurrent request merging consolidates and persists multiple write-ahead-logging operations together, enhancing I/O efficiency. In contrast, JuiceFS and Lustre rely on expensive distributed transactions to update both the file and directory metadata. CephFS do not use distributed transactions but logs writes to remote OSDs —- both of which incur significant overhead.
For getattr, FalconFS achieves 0.52–0.93$\times$ speedup over Lustre. The performance gain comes from that (a) concurrent request merging boosts server concurrency and reduce request dispatching overhead, and (b) FalconFS ’s stateless-client architecture eliminates the need for acquiring cache coherence locks (e.g., CephFS’s capability and Lustre’s intent locks).
@@
Figure 11 presents the latency of metadata operations across different DFSs. We deployed four metadata servers with a single client thread issuing requests.
Although FalconFS demonstrates superior throughput compared to other DFSs, its latency is higher than Lustre’s. This trade-off occurs because FalconFS employs concurrent request merging to batch operations, optimizing throughput at the cost of increased latency. Nevertheless, FalconFS maintains lower latency than both CephFS and JuiceFS, whose heavier software stacks introduce greater overhead.
@@ 6.3 Data Performance
In this section, we evaluate the performance of small-file access. We deploy four metadata servers and 12 storage nodes, each equipped with one NVMe SSD. We saturate the DFSs using 2560 client threads distributed across 10 client nodes. Each thread accesses 1024 pre-created files within its own private directory. To access a file, a client first opens it with the O_DIRECT flag, reads or writes all data, and then closes the file. We vary the file size from 4 KiB to 1 MiB and report the normalized throughput in Figure 12.
When the file size is smaller than 256 KiB, the throughput increases proportionally with the file size, indicating that the metadata operation IOPS is the bottleneck. When the file size is larger than 256 KiB, CephFS, Lustre and FalconFS’s throughput hits the SSD bandwidth bottleneck of 43 GiB/s for read and 16 GiB/s for write. Thanks to FalconFS’s higher metadata performance, it outperforms other DFSs in small-file access. For files no larger than 64 KiB, FalconFS achieves 7.35–21.23$\times$ speedup over CephFS, 0.86–24.87$\times$ over JuiceFS and 1.12–1.85$\times$ over Lustre. JuiceFS does not perform well in small file access due to the inefficiency of the data storage storage.

### v2

@@ 4.2.2 Statistical Load Balancing
When the coordinator detects a load imbalance using statistics reported by MNodes, it conducts the load balancing algorithm as follows.
@@
    Redistribute $F$. The coordinator selects from the two redirection methods to approach load balancing. If we use path-walk redirection, assuming it evenly distributes all files with filename $F$ to all nodes, the inode numbers of $N_{max}$ and $N_{min}$ will be $\langle{N_{max}}\rangle-\frac{n-1}{n}|F|$ and $\langle{N_{min}}\rangle+\frac{1}{n}|F|$, respectively, where $|F|$ represents the number of files named $F$ in node $N_{max}$. If using overriding redirection for $F$, we will transfer all $|F|$ files from $N_{max}$ to $N_{min}$, yielding the inode numbers of $\langle{N_{max}}\rangle-|F|$ and $\langle{N_{min}}\rangle+|F|$, respectively. The coordinator chooses the method that minimizes the maximum inode count. After the method is chosen, the redirection entry is inserted into the exception table, and the corresponding files are migrated among nodes. To ensure metadata consistency, access to the corresponding inodes is temporarily blocked during the migration.
@@ 4.3 Lazy Namespace Replication
Figure 7: Namespace Synchronization. Blue arrows represent remote lookup and red arrows represent invalidation.
Lazy namespace replication is another key enabler of the stateless-client architecture. FalconFS maintains a consistent but not necessarily complete namespace replica on each MNode and the coordinator, enabling local path resolution. To reduce the overhead of maintaining consistency across all replicas, modifications to the namespace are lazily synchronized, and an invalidation-based mechanism is adopted, inspired by cache coherence protocols and Hermes [20]. The design is guided by two principles:
@@
We will then introduce how the namespace replicas are maintained during related operations, including creating/removing a directory, changing the permissions, and renaming.
@@
Fig. 7 shows an example of how to create the directory /b. The client calculates the location of /b via the hybrid metadata indexing, and then sends a mkdir request to the corresponding MNode, i.e., $MNode_{0}$. Upon receiving the request, $MNode_{0}$ resolves the path by querying its local namespace replica to check the path existence and permissions. It also checks its inode table to confirm /b does not already exist. Once all checks are passed, $MNode_{0}$ creates an inode for /b in the inode table, adds a dentry for /b to its local namespace replica, and responds to the client.
Note that to retain the efficient single-hop processing of mkdir, $MNode_{0}$ does not proactively broadcast the new dentry to namespace replicas on other MNodes. Instead, the dentry information is fetched by other namespace replicas on demand. Specifically, when an MNode queries its local namespace replica for path resolution and finds a missing dentry, it fetches the information from the owner MNode calculated via the hybrid metadata indexing.
Fig. 7 shows an example. When $MNode_{1}$ resolves the path /b/c, it finds the dentry for /b missing in its local namespace replica. It then sends a lookup request to $MNode_{0}$ (i.e., the owner MNode) to fetch the missing dentry to complete the path resolution and continues subsequent processing.
@@
In the example shown in Fig. 7, the client sends an rmdir(/b) request to the coordinator. The coordinator acquires shared locks on all ancestor directories and an exclusive lock on the target directory to ensure path validity during execution. It then forwards the request to the directory inode’s owner, i.e., $MNode_{1}$ in the figure. $MNode_{1}$ locks /b’s inode to block subsequent lookup requests and broadcasts an invalidation request to other MNodes. Upon receipt, each MNode invalidates its local dentry of the target directory if it exists in the local namespace replica. Each MNode searches its inode table for entries whose key’s pid equals /b’s ID (i.e., /b’s children), and responds to $MNode_{1}$ the existence of any children. $MNode_{1}$ aggregates these responses. If /b has no children, $MNode_{1}$ deletes the inode and notifies the coordinator, which will release locks and respond to the client. Otherwise, $MNode_{1}$ returns -ENOTEMPTY to abort the rmdir.
@@
FalconFS leverages namespace replicas to coordinate concurrent requests. When a server (the coordinator or an MNode) processes a request, it first resolves the path component by component in its local namespace replica. Dentry locks are acquired during the resolution. To improve parallelism, the coordinator acquires shared locks on intermediate dentries and an exclusive lock on the last component, while MNodes acquire shared locks on all dentries. If a dentry is missing in the local namespace replica, the server locks the dentry after it is retrieved from its owner MNode. Thus, concurrent requests on the same server are serialized by these locks. Concurrent requests on two different MNodes will not incur a data race and thus can be executed in parallel. The last case is a request being processed on the coordinator ($Req_{C}$) and a request on an MNode ($Req_{M}$).
@@
The second case is that the invalidation is processed before $Req_{M}$ locks /a/b. In this case, $MNode_{M}$ will find the /a/b dentry invalid during path resolution of $Req_{M}$. $MNode_{M}$ then sends a lookup request to retrieve the dentry from $MNode_{C}$, who owns the /a/b inode. On $MNode_{C}$, the lookup will acquire a shared lock of the /a/b inode, which has already been locked by $Req_{C}$. Thus, $Req_{M}$ will be blocked until $Req_{C}$ completes, forming a correct serialization of the two requests. Note that when processing the invalidation request, $MNode_{M}$ will discard all lookup responses whose requests are issued before the invalidation is received.
@@ Discussion.
Lazy namespace replication introduces one remote lookup for access to a non-existent path. However, in DL workloads, such negative access is rare.
@@ 4.4 Concurrent Request Merging
Fig. 8 illustrates the metadata servers’ request-handling mechanism, which employs concurrent request merging. Each MNode initializes a fixed number of database worker threads to serve as the backend for metadata storage and prepares a connection pool. The connection pool accepts incoming requests and puts them into request queues according to the request type. An idle worker thread retrieves a queue and executes all requests in the queue in a single database transaction, with the following optimizations.
@@
Due to the tree-structured nature of the file system namespace, requests can share common near-root path prefixes. As the request queue accumulates multiple operations, the worker coalesces shared path prefixes and eliminates redundant lock acquisitions. In Fig. 8, the three create operations each walk two directories and one file. Rather than acquiring nine locks separately, the worker eliminates redundant lock acquisitions and acquires only six locks instead.
@@
Operations such as mkdir, create, and close modify the inode table. To maintain file system metadata consistency, DFSs typically warp each operation into a separate transaction to persist in atomic [34, 47, 28]. When a transaction commits, it synchronously appends the write-ahead log, leading to small writes that are unfriendly for storage. In FalconFS, as concurrent operations are batched into a single transaction, the worker coalesces small log appends into larger ones, improving the storage efficiency.
4.5 Reliability and Reconfiguration
In this section, we discuss how FalconFS supports crash consistency, high availability, and system reconfiguration.
Crash consistency.
FalconFS adopts write-ahead logging (WAL) to ensure crash consistency and atomicity of operations. Any persistent updates to the MNodes are first recorded in the WAL before being applied and visible. If an MNode fails during path resolution, update, or migration, uncommitted operations will be rolled back, and committed operations will be recovered from the log. We leverage PostgreSQL’s WAL mechanism to support single-node transactions and build a customized two-phase commit protocol upon it to ensure the consistency of operations that span multiple MNodes. Coordinator failure is treated the same as an MNode failure.
High availability.
For the high availability of the metadata service, FalconFS supports majority-based replication for MNodes and the coordinator. Each MNode and the coordinator have multiple replicas, with one primary replica serving requests and multiple secondary replicas synchronizing the primary’s state. The state synchronization is achieved by using PostgreSQL’s physical streaming replication mechanism to ship the primary’s WAL to the secondaries continuously. Once a primary replica becomes unavailable, FalconFS elects a secondary replica with the longest WAL as the new primary. FalconFS is available as long as a majority of each MNode’s replicas are available. Note that the majority-based replication is orthogonal to lazy namespace replication in §4.3.
Cluster reconfiguration.
FalconFS adopts consistent hashing to compute inode location and supports cluster reconfiguration (i.e., MNode joining and leaving) accordingly. Once the cluster needs to be resized, FalconFS migrates involved inodes to/from the added/removed MNodes. During migration, FalconFS stops serving requests. FalconFS does not support live migration since it introduces extra overhead for checking migration status on the critical path of accessing inodes.
@@ 5 Implementation of VFS Compatibility
We observe that the existing semantics provided to the lookup() method are sufficient to distinguish lookup intentions. Since Linux kernel 5.7, the VFS sets the global state flag LOOKUP_PARENT during path walk to indicate that the final component has not yet been reached — a feature designed initially for the kernel audit subsystem [6], and the flag is passed to the lookup() method. If the flag is set, the client module knows that the lookup is for an intermediate directory and returns fake attributes (e.g., $mode=0777$, along with special uiduid and gidgid values) to pass VFS checking.
@@
Fig. 9 illustrates an example of VFS shortcut. During a getattr operation for the path /a/b, the VFS resolves each component sequentially. It first looks up /, which results in a dcache hit. The VFS invokes the d_revalidate() method to validate the entry, receiving a positive response. Then, the VFS looks up /a, which misses the dcache. The lookup() method is called with the LOOKUP_PARENT flag set, returning a fake attribute. Finally, the VFS looks up /a/b, which also misses the dcache. Here, the lookup() method is invoked without flags. The module then issues a remote lookup request with the full path (/a/b) to the MNode, which executes the real path checking, executes the lookup operation, and returns the result to the module. The module returns the result to the VFS, completing the getattr operation.
@@ 6 Evaluation
Figure 11: Average latency of metadata operations.
Figure 12: Scalability with regard to concurrent clients.
@@
    FalconFS is robust under adverse conditions like client memory limitations (§6.4) and load skewness (§6.5).
@@
    The performance in real-world DL workloads (§6.8).
@@ 6.1 Environment Setup
Testbed. We conduct experiments on a cluster of 13 dual-socket machines (configurations detailed in Tab. 2). To expand the test scale, we abstract each machine into two independent nodes, with each node bound to one socket, one SSD, and one NIC, scaling the testbed to 26 nodes. We restrict server resources to 4 cores per node to ensure clients can saturate the servers’ capabilities.
@@
We compare FalconFS with CephFS 12.2.13 [49], JuiceFS 1.2.1 [3] and Lustre 2.15.6 [4]. CephFS is a widely deployed DFS in data centers. JuiceFS is an open-source DFS targeting AI and data analytics workloads, and we deploy it with TiKV 1.16.1 [8] as its metadata engine and data storage. Lustre is a high-performance DFS widely used in HPC and data centers. CephFS is accessed via the libcephfs library due to observed instability in performance when using a VFS mount point. JuiceFS and Lustre are accessed through VFS mount points, as is FalconFS unless stated otherwise. FalconFS utilizes a modified FUSE kernel module and library that incorporates the optimizations described in §5. All DFSs disable metadata and data replication.
@@ 6.2 Metadata Performance
We scale the number of metadata servers from 4 to 16 and measure the peak throughput achievable by each file system. To saturate the metadata servers’ capacity, we gradually increase the number of client threads until the throughput no longer increases. When mounting with FUSE clients, we observe that 13 client nodes are insufficient to saturate FalconFS due to bottlenecks in FUSE, so we present FalconFS’s performance using the LibFS interface, enabling each client node to generate higher concurrency. We ensure that the FUSE client and the LibFS client generate identical requests to the metadata servers; thus, given sufficient client nodes, FUSE clients would achieve comparable performance. In production environments, client nodes typically outnumber metadata servers and can fully saturate them. Fig. 10 shows the results.
[…]
@@
Fig. 11 presents the latency of metadata operations across different DFSs. We deployed four metadata servers with a single client thread issuing requests.
While FalconFS demonstrates superior throughput compared to other DFSs, its latency is higher than Lustre’s. This trade-off occurs because FalconFS employs concurrent request merging to batch operations, optimizing throughput at the cost of increased latency. Besides, FalconFS shows high p99 latency for rmdir for broadcasting invalidation requests to all MNodes and waiting for the slowest response. Nevertheless, FalconFS’s latency is comparable to CephFS and better than JuiceFS for operations other than rmdir.
Scalability with regard to concurrent clients.
Fig. 12 presents the throughput of metadata operations with an increasing number of client threads, using four metadata servers. Due to space limitations, we only present the results of create and getattr. FalconFS’s throughput scales well for both operations. When the client number is no more than 256, FalconFS’s throughput is lower than Lustre’s due to higher latency. However, as the number of clients continues to increase, Lustre’s performance saturates and FalconFS outperforms it. FalconFS’s good scalability over client number comes from that (a) the connection pool allows serving a large number of connections with a few threads, and (b) the concurrent request merging efficiently batches request executions.
@@ 6.3 Data Performance
In this section, we evaluate the performance of small-file access. We deploy four metadata servers and twelve data nodes, each equipped with one NVMe SSD. We saturate the DFSs using 2560 client threads distributed across 10 client nodes. Each thread accesses 1024 pre-created files within its own private directory. To access a file, a client first opens it with the O_DIRECT flag, reads or writes all data, and then closes the file. We vary the file size from 4 KiB to 1 MiB and report the normalized throughput in Fig. 13.
[…]